OpenStudy (anonymous):

Prove Cassini's identity using the matrix representation of the Fibonacci sequence

OpenStudy (anonymous):

oh is that all? easy! one question, whats cassini's identity? To google!

OpenStudy (anonymous):

Heh, I can open questions all over the shop now.....:-)

OpenStudy (anonymous):

http://planetmath.org/ProofofCassinisIdentity.html come across this which answeres your question, but certainly doesnt answer mine.

OpenStudy (anonymous):

out of interest. whats the formula for the fibonacci sequence

OpenStudy (anonymous):

Where's the matrix?

OpenStudy (anonymous):

\[F_n = F_{n-2} + F_{n-1}\] ?

OpenStudy (anonymous):

oh sorry yes your right there is none there

OpenStudy (anonymous):

You can represent the Fibonacci sequence as: \[\left[\begin{matrix}F_{n+1} & F_n \\ F_n &F_{n-1}\end{matrix}\right] = \left[\begin{matrix}1 & 1 \\ 1 & 0\end{matrix}\right]^n\]

OpenStudy (anonymous):

This is fairly easy to prove by induction

OpenStudy (anonymous):

define F_0 as 0, F_1 as 1, and F_2 as 1

OpenStudy (anonymous):

If the first equation is true, then\[\left[\begin{matrix}1 & 1 \\ 1 & 0\end{matrix}\right]^{n+1}=\left[\begin{matrix}1 & 1 \\ 1 & 0\end{matrix}\right]^n*\left[\begin{matrix}1 & 1 \\ 1 & 0\end{matrix}\right]\]: \[= \left[\begin{matrix}F_{n+1} & F_n \\ F_n & F_{n-1}\end{matrix}\right]*\left[\begin{matrix}1 & 1 \\ 1 & 0\end{matrix}\right]\] \[= \left[\begin{matrix}F_{n+1}+F_n & F_{n+1} \\ F_n+F_{n-1} & F_n\end{matrix}\right]=\left[\begin{matrix}F_{n+2} & F_{n+1} \\ F_{n+1} & F_n\end{matrix}\right]\]

OpenStudy (anonymous):

So: \[\left[\begin{matrix}F_{n+1} & F_n \\ F_n & F_{n-1}\end{matrix}\right] = \left[\begin{matrix}1 & 1 \\ 1 & 0\end{matrix}\right]^n \forall n \in Z_+\]

OpenStudy (anonymous):

now, take the determinant of both sides: \[F_{n+1}F_{n-1}-F_n^2=\det \left[\begin{matrix}1 & 1 \\ 1 & 0\end{matrix}\right]^n\]

OpenStudy (anonymous):

The determinant of the products of square matrices is the products of the determinants of the matrices: \[= (\det \left[\begin{matrix}1 & 1 \\ 1 & 0\end{matrix}\right])^n\]

OpenStudy (anonymous):

=\[(0-1)^n=-1^n\]